# GRASP: A Grid-Based Benchmark for Evaluating Commonsense Spatial Reasoning

### v1

@@ Abstract
Spatial reasoning, an important faculty of human cognition with many practical applications, is one of the core commonsense skills that is not purely language-based and, for satisfying (as opposed to optimal) solutions, requires some minimum degree of planning. Existing benchmarks of Commonsense Spatial Reasoning (CSR) tend to evaluate how Large Language Models (LLMs) interpret text-based spatial descriptions rather than directly evaluate a plan produced by the LLM in response to a spatial reasoning scenario. In this paper, we construct a large-scale benchmark called GRASP, which consists of 16,000 grid-based environments where the agent is tasked with an energy collection problem. These environments include 100 grid instances instantiated using each of the 160 different grid settings, involving five different energy distributions, two modes of agent starting position, and two distinct obstacle configurations, as well as three kinds of agent constraints. Using GRASP, we compare classic baseline approaches, such as random walk and greedy search methods, with advanced LLMs like GPT-3.5-Turbo and GPT-4o. The experimental results indicate that even these advanced LLMs struggle to consistently achieve satisfactory solutions.
1 Introduction
@@
The ability to use interpreted spatial information to navigate and interact with the environment in a commonsensical manner is termed Commonsense Spatial Reasoning (CSR) ability. It is essential to effectively navigate around environments. For example, an agent must understand the fundamental spatial principles such as ‘if an obstacle is in the way, an alternative path that goes around the obstacle should be considered’ and ‘if two resources are close together, it is more efficient to collect them in one trip rather than multiple trips.’ Such CSR enables the agent to make effective decisions with regard to the spatial layout, such as prioritizing the collection of nearby resources before exploring further and adapting to obstacles by rerouting effectively.
@@
GRASP is a set of environments that consists of 2000 different grid instances and 8 distinct combinations of agent constraints, resulting in 16000 environments. Each grid is constructed as a two-dimensional array of cells, each of which can be empty, an obstacle, or contain a unit of energy. The objective for the agent is to collect as much energy as possible within a fixed number of steps, simulating real-world scenarios where agents must navigate through the environments and plan a path to collect resources. The environment includes various patterns of energy arrangement, such as random, vertically-skewed, horizontally-skewed, clustered, and spiral distributions. These patterns challenge the agent’s ability to recognize the differences and adapt its strategy based on different distributions of resources. Additionally, obstacles are randomly placed within the grid and the agent’s starting position is randomized within specified inner or outer regions of the grid, testing the agent’s adaptability to different starting scenarios. Furthermore, we introduce various constraints that affect the behaviors of the agent, starting with a limited set of actions that include basic movements and, in a more complex setup, additional diagonal movements. Additional Constraints, such as the maximum number of energy units the agent can carry and the cost associated with each step, are designed to further test the agent’s CSR and associated planning abilities.
2 Related Work
[…]
Additionally, BabyAI [4] provides a grid-based platform to support the research of including humans in the loop for grounded language learning. Minigrid [5] provides modular and customizable grid-based environments designed for goal-oriented tasks in the reinforcement learning paradigm. However, these benchmarks primarily focus on evaluating an agent’s ability to follow natural language instructions with spatial information, while our benchmark investigates the agent’s ability to autonomously navigate the grid and collect resources using spatial information. More importantly, they see the task as needing further model optimization, while we argue that spatial reasoning should be treated as a commonsense ability and evaluated accordingly [20]. Moreover, in our grid environment, the agent begins with complete knowledge of the grid’s layout and all necessary task information. Unlike other benchmarks that require agents to explore and actively gather data, our framework provides all information upfront, eliminating the need for sequential interaction with the environment. Table 1 provides a comparison between GRASP and other related benchmarks.
Table 1: GRASP in Comparison to Other Benchmarks
Benchmark Spatial Representation Planning Required? Optimization Needed? Information Scope Spatial Commonsense Needed?
Minigrid [5] visual yes yes partially no
BabyAI [4] visual yes yes partially no
PPNL [2] textual yes yes fully no
SPARTQA [16] textual no no fully no
StepGame [22] textual no no fully no
GRASP (ours) hybrid yes no fully yes
3 GRASP: Benchmark Construction
@@
GRASP evaluates the commonsense spatial reasoning ability of LLMs by testing their capability to navigate in a grid-base environment while assigned an energy collection task. The benchmark is a combination of grid environments and agent constraints. The grid environment, $G$, can be understood as a 2D square array of cells $C_{ij}$, indexed from 0 to $k$ and from left to right and top to bottom, with $i$ representing row index and $j$ representing column index: $C_{00},...,C_{0k},C_{10},...,C_{1k},...,C_{k0},...,C_{kk}$. A cell can either be empty ($C_{ij}=Null$), contain one unit of energy ($C_{ij}=E$), have an obstacle ($C_{ij}=O$), or be the starting point of the agent ($C_{ij}=A$).
@@
As for the agent constraints, the agent is allowed pre-defined ($n$) steps. At each step, the agent can choose from a set of movement-related actions ($\mu$) that allows it to move around the grid environment and a set of resource-related actions ($\rho=\{TAKE,DROP\}$) that enables it to take the energy from or drop the collected energy to the cell that it is in. The agent’s objective is to collect as much energy as possible in $n$ steps. In section 3.1, we introduce three different controls that can affect the layout of the grid environment. In section 3.2, we discuss three types of constraints that the agent must follow in accomplishing the task in our grid environment. In 3.3, we explain how such a grid environment is represented to an LLM-based agent. Figure 1 shows the entire pipeline for constructing one instance of the grid environment.
3.1 Grid Environments
@@ 3.2 Agent Constraints
3.3 Grid Representation to An LLM-based Agent
@@
4 Experiments
@@
Two baseline approaches are evaluated. The first is the random walk method. This method simulates an agent who has no knowledge about the grid, the agent constraints regarding the limit on the number of energy that the agent can carry, and the cost of each action, but is only aware of the set of available actions. The random walk method randomly selects one out of the four or eight movement-related actions, depending on the available actions. Following each of the movement-related actions, this approach will select the TAKE action to take the energy from the cell, whether or not the cell has any energy. This ‘move and take’ behavior will continue for 6 times. Then, this approach will select the complement of the movement-related actions that were selected before in reverse order to get back to the starting position and finally select DROP to drop all the energy. The complement of a movement-related action is defined as the complementary action that exactly reverses the movement. For example, if the initial movement-related action was ‘UP,’ the complementary action would be ‘DOWN.’ In total, this will take 19 steps, which is within the required 20 steps.

### v2

@@ Abstract
Spatial reasoning, an important faculty of human cognition with many practical applications, is one of the core commonsense skills that is not purely language-based and, for satisfying (as opposed to optimal) solutions, requires some minimum degree of planning. Existing benchmarks of Commonsense Spatial Reasoning (CSR) tend to evaluate how Large Language Models (LLMs) interpret text-based spatial descriptions rather than directly evaluate a plan produced by the LLM in response to a specific spatial reasoning problem. In this paper, we construct a large-scale benchmark called GRASP, which consists of 16,000 grid-based environments where the agent is tasked with an energy collection problem. These environments include 100 grid instances instantiated using each of the 160 different grid settings, involving five different energy distributions, two modes of agent starting position, and two distinct obstacle configurations, as well as three kinds of agent constraints. Using GRASP, we compare classic baseline approaches, such as random walk and greedy search methods, with advanced LLMs like GPT-3.5-Turbo, GPT-4o, and GPT-o1-mini. The experimental results indicate that even these advanced LLMs struggle to consistently achieve satisfactory solutions.
@@
The ability to use interpreted spatial information to navigate and interact with the environment in a commonsensical manner is termed Commonsense Spatial Reasoning (CSR). It is essential to effectively navigate around environments. For example, an agent must understand the fundamental spatial principles such as ‘if an obstacle is in the way, an alternative path that goes around the obstacle should be considered’ and ‘if two resources are close together, it is more efficient to collect them in one trip rather than multiple trips.’ Such CSR enables the agent to make effective decisions with regard to the spatial layout, such as prioritizing the collection of nearby resources before exploring further and adapting to obstacles by rerouting effectively.
@@
GRASP is a set of environments that consists of 2,000 different grid instances and eight distinct combinations of agent constraints, resulting in 16,000 environments. Each grid is constructed as a two-dimensional array of cells, each of which can be empty, has an obstacle, or contains a unit of energy. The objective for the agent is to collect as much energy as possible within a fixed number of steps, simulating real-world scenarios where agents must navigate through the environments and plan a path to collect resources. The environment includes various patterns of energy arrangement, such as random, vertically-skewed, horizontally-skewed, clustered, and spiral distributions. These patterns challenge the agent’s ability to recognize the differences and adapt its strategy based on different distributions of resources. Additionally, obstacles are randomly placed within the grid, and the agent’s starting position is randomized within specified inner or outer regions of the grid, testing its adaptability to different starting scenarios. Finally, we introduce various constraints that affect the behaviors of the agent, starting with a limited set of actions that include basic movements and, in a more complex setup, additional diagonal movements. Additional constraints, such as the maximum number of energy units the agent can carry and the cost associated with each step, are designed to further test the agent’s CSR and associated planning abilities.
We use two baseline agents to set a foundation for evaluating the performance of LLM agents in the GRASP benchmark. Experimental results reveal that GPT-3.5-Turbo, GPT-4o, and GPT-o1-mini all underperform compared to the first baseline method, which is based on greedy search. We also find that GPT-3.5-Turbo falls short of the second, simpler random walk baseline. Among the LLM-based agents, GPT-o1-mini exhibits the most promising performance, achieving, on average, the best results and demonstrating reasonable adaptability to constraints such as obstacles and limited allowed movements. However, the introduction of constraints such as energy costs per step or energy-carrying limit significantly impacts all LLM-based agents’ performance, showing that there are still significant challenges to deploying them on reasoning and planning problems that resemble CSR. Our findings ultimately support GRASP’s potential to benchmark and reveal important differences in agents’ CSR abilities and offer valuable insights into the capability and shortcomings of LLM-based agents in practical spatial reasoning tasks.
Results
Benchmark Overview
@@ Results
We designed GRASP specifically to benchmark LLM-based agents’ CSR abilities. GRASP tests LLM-based agents’ capability to navigate in a grid-based environment, represented using simple text characters, while assigned an energy collection task. The goal of the agent is to collect as much energy as possible within a certain number of steps. Figure 1 shows the entire pipeline for constructing one instance of the grid environment. The benchmark is a combination of grid environments and agent constraints. For one grid environment, there can be five different types of energy distribution: random, vertically-skewed, horizontally-skewed, cluster, and spiral. These five types are used to create a broad enough set of scenarios that cannot be ‘gamed’ by agents through memorization or obvious heuristics. Grid environments can also contain obstacles that block specific grid cells, making those cells inaccessible. The obstacles are randomly placed inside the grids to introduce challenges and encourage strategic navigation. The agent’s starting position can be either the inner position or the outer position. The inner position refers to cells near the center of the grid, while the outer position is the cells near the edges. These starting positions allow the benchmark to evaluate the agent’s performance under different spatial conditions.
There are three kinds of agent constraints:
  1. 1.
    First, the agent is allowed two types of movement-related actions: one with four directions and the other with eight directions. These movement options assess the agent’s ability to navigate the grid with varying levels of freedom.
  2. 2.
    Second, the agent can have an energy carry limit, which restricts the amount of energy the agent can hold at any given time. This constraint encourages the agent to plan its energy collection and usage strategically.
  3. 3.
    Finally, the agent can incur an energy cost per step, which determines the amount of energy consumed with each step taken. This cost introduces a trade-off between exploration and resource conservation, challenging the agent to optimize its movements. The agent is also restricted to only 20 steps. The detailed parameters and processes are described in Methods.
Experimental Setup and Parameters
We start by briefly describing the parameters of the experimental study for evaluating CSR in LLM-based agents using GRASP.
Baselines: To evaluate agent performance on GRASP, we begin by establishing two baseline methods that represent different levels of capabilities. The baselines establish reference points for interpreting the results of the LLM-based agents. The first baseline is the random walk method. It represents a lower-bound performance and serves as a control to assess the inherent difficulty of the task. It simulates an agent who has no knowledge about the grid, the agent constraints regarding the limit on the number of energy that the agent can carry, and the cost of each action, but is only aware of the set of available actions. An agent who employs this approach randomly selects actions from the set of available movements. Basically, this baseline represents how a completely uninformed agent would perform in GRASP, offering a minimum performance level against which more sophisticated agents can be compared.
The second baseline is the greedy search method, which continues to be popular in many practical planning scenarios[29, 30, 31] where devising optimal solutions (e.g., by solving an elaborately specified integer linear program) is combinatorially explosive. This method simulates an agent with perfect knowledge of the grid environment but no planning abilities and no knowledge about the agent constraints except the available movement-related actions. The greedy agent always selects the action that maximizes immediate energy gain. This baseline demonstrates how full environmental knowledge can enhance performance. Implementation details are provided in Methods.
LLM-based agents: In addition to the baseline approaches, we evaluate the performance of agents based on three LLMs: GPT-3.5-Turbo (i.e., gpt-3.5-turbo-0125), GPT-4o, and GPT-o1-mini. All three models are commercial, with GPT-4o and GPT-o1-mini being close to the state-of-the-art at the time of writing. The models were chosen due to their varying levels of capability and costs. GPT-3.5-Turbo is included as a cost-effective, widely accessible model, while GPT-4o is evaluated as a more advanced model known for its superior reasoning capabilities. GPT-o1-mini is both more expensive and computationally stronger compared to GPT-3.5-Turbo and GPT-4o, making it an interesting point of comparison. The first two models are used with a temperature setting of 0 to ensure reproducibility. However, because GPT-o1-mini does not allow modification to its default temperature, it is kept at 1.
These models are tested using zero-shot prompting . To construct the prompt used for the LLMs, we utilize all available information, both from the grid environment and agent constraints. Instructions and agent constraints are provided to LLMs in the system prompt, and the grid is provided in the user prompt; the LLMs are required to provide a full list of actions in the response, with no further interaction with the environment. GPT-o1-mini can only accept user prompts, so all information is provided in the user prompt. The specific prompts used are detailed in Methods. The experiments are conducted using one-tenth of the entire GRASP benchmark, for grid instances indexed from 0 to 9, except for GPT-o1-mini, which is evaluated using grid instances indexed from 0 to 4.
Performance Evaluation
$\downarrow$ Control &\&& its Values $\searrow$ Agent $\rightarrow$ Random Walk Greedy Search GPT-3.5- Turbo GPT-4o GPT-o1-mini
Length Energy Length Energy Length Energy Length Energy Length Energy
Energy Distribution Random 19.0 -0.80 18.8 0.37 19.7 -1.96 16.6 -0.99 18.5 -0.44
Vertically-skewed 19.0 -0.98 18.7 0.33 19.6 -1.66 17.4 -1.40 18.6 -0.30
Horizontally-skewed 19.0 -0.97 18.6 0.47 19.7 -2.11 17.3 -1.26 18.4 -0.32
Cluster 19.0 -1.60 18.8 0.37 18.6 -2.07 16.7 -1.06 18.2 -0.33
Spiral 19.0 -1.38 18.7 0.20 19.4 -2.32 17.1 -1.34 18.3 -0.17
Obstacle Yes 19.0 -1.18 18.7 0.31 19.3 -2.08 17.2 -1.45 18.3 -0.32
No 19.0 -1.11 18.7 0.39 19.5 -1.97 16.8 -0.96 18.6 -0.31
Starting Position Inner Position 19.0 -1.06 18.7 0.46 19.4 -1.90 17.0 -1.17 18.7 -0.35
Outer Position 19.0 -1.23 18.7 0.23 19.4 -2.14 17.0 -1.25 18.3 -0.27
Movement-related Action Set $\mu_{1}$ (4 directions) 19.0 -1.21 18.5 0.80 19.1 -1.89 16.8 -1.13 18.5 -0.19
$\mu_{2}$ (8 directions) 19.0 -1.08 18.9 -0.10 19.7 -2.15 17.3 -1.28 18.5 -0.43
Energy Carrying Limit No Limit 19.0 -0.89 18.7 1.50 19.8 -2.00 16.7 -0.85 18.3 0.24
2 Units 19.0 -1.40 18.7 -0.81 19.0 -2.05 17.3 -1.57 18.6 -0.87
Energy Cost Per Step 0 Unit 19.0 1.68 18.7 3.14 19.4 1.25 17.2 1.34 18.4 2.39
0.3 Unit 19.0 -3.97 18.7 -2.44 19.4 -5.30 16.9 -3.76 18.6 -3.02
Average 19.0 -1.14 18.7 0.35 19.4 -2.02 17.0 -1.21 18.5 -0.31
Table 1 displays the average number of steps taken and units of energy collected and dropped for the two baseline agents (random walk and greedy search) and the three LLM-based agents (GPT-3.5-Turbo, GPT-4o, and GPT-o1-mini), evaluated using GRASP. Most Energy results are negative numbers because they are averaged across all control settings, including a 0.3 unit energy cost per step that can lead to a negative energy result if the energy cost exceeds the energy collected. When considering the overall performance of each agent, we find that the greedy search agent performs the best across all settings, while GPT-3.5-Turbo performs the worst in 14 out of 14 settings. Additionally, GPT-4o performs better than GPT-3.5-Turbo, indicated by an unpaired t-test ($t(3198)=6.78$, $p<0.001$), but worse than GPT-o1-mini ($t(2398)=-6.27$, $p<0.001$) and is roughly on par with the random walk agent ($t(17598)=-0.48$, $p=0.632$). All LLM-based agents perform worse than the greedy search agent ($p<0.001$), with only GPT-o1-mini performing better than the random walk method ($t(16798)=-7.64$, $p<0.001$). On the length of the steps, because the path of the random walk method is mechanically determined, it maintains consistent lengths of 19. The LLM-based agents exhibit consistent step counts within each individual model across different controls. However, when comparing their average length, GPT-4o consistently achieves the shortest path lengths, often between 16 to 17 steps, followed by GPT-0.1-mini, which has slightly longer paths, taking between 18 to 19 steps. In contrast, GPT-3.5-Turbo consistently produces the longest paths among the LLM-based agents, with about 19 to 20 steps.
Focusing on the performance within each control, we found that GPT-3.5-Turbo achieves the best Energy results under the horizontally-skewed energy distribution, while GPT-4o performs the best when the energy distribution is random. Interestingly, GPT-o1-mini got the highest Energy result under the spiral distribution. However, through paired t-tests, we found that the differences are not statistically significant ($p>0.01$). Additionally, we observed Energy differences of 0.49 and 0.72 between the two settings of ‘Obstacle,’ and ‘Energy Carrying Limit’ for GPT-4o. This agent performs best without obstacles ($t(1598)=3.03$, $p<0.01$) and without any energy carrying limit ($t(1598)=4.51$, $p<0.001$). For GPT-o1-mini, the difference between the two settings of ‘Energy Carrying Limit’ is 1.11, suggesting that this agent performs best without any energy carrying limit ($t(798)=4.62$, $p<0.001$). However, these differences are less apparent and not significant for GPT-3.5-Turbo, with the maximum difference being only 0.26 units of energy in the ‘Movement-related Action Set’ control ($t(1598)=1.45$, $p=0.148$).
All three LLM-based agents achieve notably lesser energy when given the constraint regarding the energy cost per step than when given no such constraint, with the minimum difference being 5.1 units of energy from GPT-4o ($p<0.001$ for all LLM-based agents). Additionally, the “Movement-related Action Set’ and ‘Starting Position’ controls seem to play a lesser role in affecting the performance of GPT-3.5-Turbo, GPT-4o, and GPT-o1-mini, as the difference in the amount of energy collected and dropped is less apparent (with the maximum difference being less than 0.26 and $p>0.1$ for all LLM-based agents).
Figure 2: The path formed by the actions of the five agents (Random Walk, Greedy Search, GPT-3.5-Turbo, GPT-4o, and GPT-o1-mini) in the shown grid environment. The agents are under the following constraints: unlimited energy carrying limit, zero energy cost per step, and the $\mu_{1}$ movement-related action set (i.e., UP, DOWN, LEFT, RIGHT). Arrows indicate the movement-related actions, and a yellow star next to each cell signifies the agents’ ‘TAKE’ action in that cell.
Figure 2 illustrates a typical action path of the five agents within the shown instance of the grid environment. We observe that the random walk agent wanders through the grid without a clear objective and gives the ‘TAKE’ action in cells without energy, while the greedy search agent exhibits a more object-oriented behavior, moving directly towards cells with energy and taking it. Focusing on the LLM-based agent, we find that GPT-3.5-Turbo includes cells with energy in its path but often fails to collect energy, repeatedly attempts to move through the cell with an obstacle, and does not return to the starting cell to drop the collected energy. In contrast, the GPT-4o agent demonstrates more intelligent behavior by actively going towards cells with energy and collecting them while avoiding obstacles, yet it still shows inefficient behaviors, such as missing nearby energy cells with energy on them and taking extra and unnecessary steps to achieve the same objectives. On the other hand, GPT-o1-mini exhibits even more sophisticated behavior in terms of both path efficiency and task completion. It prioritized collecting energy from cells in close proximity. However, we still observe a certain degree of inefficiency in that it sometimes overlooked nearby cells with energy or took longer paths than necessary to complete its objectives. These behaviors also reflect the performance of each agent observed in Table 1.
Discussion
In this paper, we introduce GRASP, a comprehensive benchmark designed to evaluate the Commonsense Spatial Reasoning (CSR) abilities of LLMs within structured grid environments. GRASP assesses agents’ abilities to navigate the grid environments and collect energy, presenting 16,000 diverse environments characterized by five types of energy distributions, two distinct obstacle configurations, two modes of agent starting position, and three kinds of agent constraints, including two sets of movement-related actions, two limits on the amount of energy that can be carried, and two different energy costs for each step. We evaluate the performance of two classic algorithms (random walk and greedy search) and three advanced LLMs (GPT-3.5-Turbo, GPT-4o, and GPT-o1-mini) using GRASP. Our findings show that GPT-3.5-Turbo under-performs compared to the random walk agent, often unable to successfully navigate the grid, avoid obstacles, and return to the starting point to drop the collected energy, suggesting a lack of CSR abilities. On the other hand, GPT-4o performs similarly to the random walk agent but is still less efficient than the greedy search agent. While GPT-4o can recognize spatial relationships and navigate around the grid environments, showing a primitive capability of CSR, its use of spatial information to come up with efficient plans remains limited. GPT-o1-mini, however, demonstrates superior performance compared to the random walk agent and other LLMs evaluated here. It is able to leverage spatial reasoning to plan efficient paths, actively prioritizing nearby energy cells and avoiding obstacles with a higher degree of accuracy. Despite its relative success, GPT-o1-mini still under performs when compared to the greedy search agent. For instance, it occasionally fails to optimize its route fully, leading to unnecessary or optimized steps.
Our work builds upon and extends prior investigations into the spatial reasoning abilities of LLMs, moving beyond existing benchmarks to introduce a more practical evaluation framework. SpartQA [22], for example, evaluates spatial reasoning using a textual Question-Answering (QA) benchmark, which is constructed using grammar and reasoning rules to automatically generate a spatial description of visual scenes and corresponding QA pairs. Similarly, StepGame [23] is another multi-hop spatial reasoning benchmark that investigates the navigation ability of LLMs through complex spatial descriptions, while Liu et al. [25] explored spatial commonsense by exploiting visual signals and evaluating LLMs’ understanding of the relative scales of objects and the positional relationship between people and objects. On the other hand, Comsa et al. [32] focused on reasoning with spatial prepositions, which is essential for accurately describing and interpreting spatial relationships. However, these benchmarks focus on interpreting and understanding spatial relationships in textual or visual formats, while GRASP evaluates the ability to utilize spatial information to accomplish related tasks. Our distinct approach assesses the practical application of spatial reasoning, requiring models to navigate and make plans based on spatial information. Another key distinction is that these benchmarks frequently depend on textual descriptions of scenes to convey spatial relationships. In contrast, our benchmark directly integrates text renderings of the environment to insert spatial information.
Additionally, platforms such as BabyAI [24] provide a grid-based platform to support the research of including humans in the loop for grounded language learning. Additionally, Minigrid [33] provides modular and customizable grid-based environments designed for goal-oriented tasks in the reinforcement learning paradigm. However, these benchmarks primarily focus on evaluating an agent’s ability to follow natural language instructions with spatial information, while our benchmark investigates the agent’s ability to autonomously navigate the grid and collect resources using spatial information. More importantly, they see the task as needing further model optimization, while we argue that spatial reasoning should be treated as a commonsense ability and evaluated accordingly [13]. Moreover, in our grid environment, the agent begins with complete knowledge of the grid’s layout and all necessary task information. Unlike other benchmarks that require agents to explore and actively gather data, our framework provides all information upfront, eliminating the need for sequential interaction with the environment.
Our study comes with several limitations. Firstly, the synthetic nature of GRASP’s grid environments may not fully capture the complexity of real-world CSR tasks, which include more dynamic and unpredictable environments with a greater variety of objects and incomplete or even hidden spatial information. Furthermore, GRASP focuses on commonsense spatial reasoning but does not consider the multi-modal nature of many real-world tasks where abundant visual information is available. Despite the current limitations and potential hallucination problems in LLMs’ multi-modal abilities, a reliable visual interpretation capability could provide the models with a richer set of spatial information than text alone. Also, we only consider scenarios with a single agent, whereas real-world environments often involve multiple agents, requiring more sophisticated behaviors, such as interacting and planning together with the other agents. Finally, our evaluation was limited to zero-shot prompting methods for LLMs, potentially underestimating their true capabilities.
To address these limitations, GRASP will be expanded in the future to include more dynamic environments that better simulate multi-modal CSR tasks, preferably developed using real-world data instead of purely synthetic ones. Additionally, with the rapid advancement of multi-modal foundational models, it is critical to integrate datasets that combine visual and textual spatial information to better capture the complexity of real-world CSR and provide a more comprehensive evaluation of the models. Extending the benchmark to scenarios involving multiple agents would also more accurately reflect the intricacies of real-world environments. Last but not least, investigating more sophisticated prompting methods could improve the performance of LLMs in these tasks, providing a more holistic evaluation of their commonsense spatial reasoning abilities.
In conclusion, GRASP is a novel benchmark for understanding and evaluating the commonsense spatial reasoning abilities of LLMs, albeit with some limitations. Our findings showcase the current limitations of even advanced LLM-based agents in efficiently completing CSR tasks and suggest the need for continued research and development in this area. As a structured and diverse benchmark, GRASP also provides a foundation for benchmarking future advancements in LLM capabilities, ultimately contributing to the development of more robust and intelligent systems capable of performing a wide range of commonsense planning tasks.
Methods
Detailed Overview of Benchmark Construction
@@ Methods
In general, the grid environment, $G$, can be understood as a 2D square array of cells $C_{ij}$, indexed from 0 to $k$ and from left to right and top to bottom, with $i$ representing row index and $j$ representing column index: $C_{00},...,C_{0k},C_{10},...,C_{1k},...,C_{k0},...,C_{kk}$. A cell can either be empty ($C_{ij}=Null$), contain one unit of energy ($C_{ij}=E$), have an obstacle ($C_{ij}=O$), or be the starting point of the agent ($C_{ij}=A$).
@@
The agent is allowed pre-defined ($n$) steps. At each step, the agent can choose from a set of movement-related actions ($\mu$) that allows it to move around the grid environment and a set of resource-related actions ($\rho=\{TAKE,DROP\}$) that enables it to take the energy from or drop the collected energy to the cell that it is in. The agent’s objective is to collect as much energy as possible in $n$ steps.
Grid Environments
@@
Grid Representation to An LLM-based Agent
@@
Implementation Details for Experimental Study
Baseline agents
@@
The random walk method randomly selects one out of the four or eight movement-related actions, depending on the available actions. Following each of the movement-related actions, this approach will select the TAKE action to take the energy from the cell, whether or not the cell has any energy. This ‘move and take’ behavior will continue 6 times. Then, this approach will select the complement of the movement-related actions that were selected before in reverse order to get back to the starting position and finally select DROP to drop all the energy. The complement of a movement-related action is defined as the complementary action that exactly reverses the movement. For example, if the initial movement-related action was ‘UP,’ the complementary action would be ‘DOWN.’ In total, this will take 19 steps, which is within the required 20 steps.